Day shift summary, 12/12/99

AG (ag@fnal.gov)
Sun, 12 Dec 1999 16:46:54 -0500 (EST)

On shift: TA, AG

0700 COSMICS run 16001 is still in progress, got some ktevana errors.

0800 TRD gas system is back on.

1350 Stan from MCR phones to tell that accelerator is doing pretty
well and might be ready within 2-3 hours. Beam can be delivered as
early as today's evening shift.

Page Sasha L., try to get hold of John S. ...

1445 Jason H. turns on ET H2O cooling in anticipation of beam and starts
new COSMICS run to blast array with laser to verify HCC bits.

1500 Got hold of John S.
-> he'll be at ktev hall soon to wrap up trigger 1 skim jobs.

1530 Jason verifies HCC/ET status (see details in KTeV Log XXIII, p.62).
He swapped board #55 in TE-8 with spare board #30 to fix a missing
bit. More problems are discovered -> Jason + Sydney (over the phone)
continue work ...

* * *

Follow up on the next shift:

Please page Ed after beam's been delivered for about half an hour.